OpenStudy (anonymous):

Help with inequalities pleasee? Three fourths of a number decreased by 9 is at least 42. Define variable, write inequality, and solve

OpenStudy (mathlegend):

Let me ask you this... 3/4 of what?

OpenStudy (mathlegend):

@SydneyKohl

OpenStudy (mathlegend):

I need feedback...

OpenStudy (texaschic101):

3/4x - 9 >= 42 (add 9 to both sides) 3/4x >= 42 + 9 3/4x >= 51 (divide both sides by 3/4) x >= 51/(3/4) x >= 51 * 4/3 x >= 204/3 x >= 68

OpenStudy (texaschic101):

do you understand or do you have questions ?